OpenStudy (anonymous):

Find all values of z such that: 2/z+5=5/z-4

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

\[\frac{2}{z}+5=\frac{5}{z}-4\] \[\frac{2}{z}-\frac{5}{z}=-4-5\] \[\frac{2-5}{z}=-9\] \[\frac{-3}{z}=-9\] \[-3=-9z\] \[z=\frac{-3}{-9}\] \[z=\frac{1}{3}\]